# Recurrence method in Non-Hermitian Systems

### v1

@@ I Introduction
Non-Hermiticity is a significant factor that can not be overlooked in both classical and quantum systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. In these systems, non-Hermiticity gives rise to various unconventional phenomena such as non-Hermitian skin effect [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], non-Bloch parity-time symmetry [30, 31, 32, 33], non-Hermitian topology [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], generalized Brillouin zone [46, 44, 47, 48, 49], and exceptional points [50, 51, 52, 53, 54, 55]. One intriguing property of non-Hermitian systems is the breakdown of spectral stability under different boundary conditions [56, 57, 58, 59, 60]. Specifically, it is widely believed that the energy levels of Hermitian systems with an open boundary condition (OBC) can be reproduced mainly by the Bloch band theory under periodic boundary conditions (PBCs), even if the translation symmetry is broken. In non-Hermitian systems, however, the energy spectra under OBC change drastically and collapse into open arcs encircled by the loop-shape periodic boundary spectra [61, 44, 21, 62]. It is widely known that these non-Hermitian OBC spectra can be obtained by the non-Bloch theory, which extends the Brillouin zone to the generalized Brillouin zone (GBZ) in non-Hermitian systems [44, 48, 47, 21]. Unfortunately, the determination of GBZ is somewhat challenging; therefore, it may be necessary to use the auxiliary GBZ technique [47], which still involves additional concepts and may not apply to complicated multi-band non-Hermitian systems, such as non-Hermitian Hofstadter butterflies [47]. Furthermore, non-Bloch theory primarily emphasizes the bulk OBC spectra relevant to the GBZ, and specialized tools for edge spectra of non-Hermitian systems, intimately connected with topological properties, are still largely lacking.
@@
In this paper, we propose an alternative method based on the recurrence relations of the eigenvalue equations in non-Hermitian systems under OBC instead of the GBZs so that we successfully avoid the limitations of non-Bloch band theory. First, we conclude that the characteristic polynomial $D_{N}(E)=\det(E-H_{N})$ of any non-Hermitian tight-binding Hamiltonian under OBC satisfies a recurrence relation with constant coefficients independent of the matrix size $N$. Then, we can represent $D_{N}(E)$ with the solutions $z_{i}(E)$ of the corresponding characteristic equation. The main result is that the OBC spectra include the energies such that (i) two maximal complex solutions $z_{i}$ and $z_{j}$ have the same modulus, or (ii) when representing $D_{N}$ by $z_{i}$, the maximal complex solution term is zero. We will argue that the above conditions are related to the bulk and edge parts of OBC spectra, respectively.
@@
As an application of the recurrence method, we derive a general expression for the bulk OBC spectra of non-Hermitian multi-band Hamiltonians with nearest-neighbor hoppings. This expression has the form of polynomial equations that allow us to analytically or numerically solve the OBC spectra without either diagonalization or determination of the GBZ. Thus, the recurrence method avoids typical accuracy problems in diagonalizing non-Hermitian matrices and significantly reduces the computational complexity. Furthermore, our formalism straightforwardly determines non-Hermitian Hamiltonians’ edge spectra, which are unavailable from the conventional non-Bloch band theory. Based on this framework, we can analyze the effects of the overall number of sites (e.g., the size-parity effect) [63, 64, 65] and the boundary perturbations [66, 67, 68, 69] on the edge spectra. For example, we study the non-Hermitian Hofstadter model in two dimensions as an interplay between the non-reciprocity and magnetic field [70]. Our results provide robust evidence for the complex semiclassical theory [71] and topological characterizations of non-Hermitian systems [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. Generally, our work offers a novel technique and perspective for non-Hermitian systems’ OBC spectra and physics.
@@ II Recurrence Method in non-Hermitian systems
where $d_{i}(E)$ are determined by the initial conditions $D_{0}(E)=1,D_{k}(E),\cdots,D_{(s-1)k}(E)$ of the recurrence relation in Eq. (3). The OBC spectrum in the thermodynamic limit $N\rightarrow\infty$ is comprised of values of $E$ satisfying either of the following two conditions [75, 74]:
@@
Note that condition (i) contains an infinite number of $E$ and a similar form with the GBZ condition $|\beta_{p}(E)|=|\beta_{p+1}(E)|$ in the non-Bloch band theory, thus corresponding to the continuous bulk OBC spectra of the non-Hermitian systems. On the other hand, condition (ii), combined with the initial conditions of Eq. (5), gives several polynomial equations about $E$ and thus a discrete set of solutions. These polynomial equations are generally quite different from the non-polynomial form of condition (i). Hence, we expect condition (ii) to contain the edge spectra of non-Hermitian systems under OBCs, which is beyond the scope of the non-Bloch band theory and the GBZ with only information of the bulk spectra [48, 47].
The recurrence method is remarkably useful for non-Hermitian models, especially models with large unit cells ($k\gg 1$) that are exceedingly complex from a non-Bloch theory perspective. As we will show, even for a large unit cell, the recurrence method gives an analytic expression for the OBC spectra of 1D non-Hermitian nearest-neighbor tight-binding models in the thermodynamic limit $N\rightarrow\infty$. Computationally, on the other hand, the recurrence method transfers the OBC spectrum problem to the problem of solving polynomials of degree $k$, which reduces the time complexity to $O(Nk^{2})$ with the Durand-Kerner algorithm [76, 77, 74].
@@ III.1 Periodic nearest-neighbor-hopping non-Hermitian models
which is an analytical expression for the OBC spectra of multiple-band Hamiltonians in Eq. (6) through the recurrence method. Therefore, our formalism transforms the OBC spectra problems of Eq. (6) to a sequence of polynomial equations in Eq. (13). For a lattice period or sublattice size $k$ that is arbitrarily large, we can solve Eq. (13) numerically through an iterative method called the Durand-Kerner method [76, 77, 74] with manageable cost $\sim O(Nk^{2})$, advantageous over the non-Bloch band theory.
@@ III.2 Non-Hermitian Rice-Mele model
where $w_{1/2}$ denote the intercell hoppings between the adjacent unit cells, $v_{1/2}$ denote the intracell hoppings between sites A and B in the same unit cell, and $\Delta$ is a staggered onsite potential, as shown in Fig. 1(a). The $V=0$ case recovers the non-Hermitian SSH model in Eq. (15) with vanishing onsite potentials. The asymmetric intracell/intercell hoppings, e.g., $v_{1}=v-\gamma$ versus $v_{2}=v+\gamma$ with a finite $\gamma$, introduce non-Hermiticity into the model and can be realized in experiments by reservoir engineering in the ultra-cold atom systems and imaginary gauge field in coupled optical micro-ring resonators [35, 82, 83].
@@
Figure 1: We compare the OBC (red curves and dots) and PBC (blue curves) spectra of the non-Hermitian Rice-Mele model in Eq. (21) with two sublattices A and B and asymmetric hoppings $v_{1}\neq v_{2}^{*}$ or $w_{1}\neq w_{2}^{*}$, illustrated in (a). The OBC spectra are fully real when both $v_{1}v_{2},w_{1}w_{2}>0$, irrespective of $V$: (b) $v_{1}=2$, $v_{2}=1$, $w_{1}=1.2$, $w_{2}=0.8$, and $V=0.5$; (c) $v_{1}=0.8$, $v_{2}=0.2$, $w_{1}=1.2$, $w_{2}=0.8$, and $V=0.5$. However, when both $v_{1}v_{2},w_{1}w_{2}<0$, a non-Bloch PT symmetry breaking may occur: (e-g) $v_{1}=-0.8$, $v_{2}=0.2$, $w_{1}=-1$, and $w_{2}=1$, thus the transition points are at $V_{c}=1.4$. The spectrum is complex for (d) $V=0.5$ and (e) $V=1.2$ at a critical point at (f) $V=1.4$ and becomes fully real for (g) $V=1.5$. Note that $|v_{1}v_{2}|>|w_{1}w_{2}|$ and thus no edge state exists in (b), while $|v_{1}v_{2}|<|w_{1}w_{2}|$ and two isolated edge states appear at $E=\pm V$ (red dots) in (c-g).
@@ IV.1 Edge spectra by recurrence method
Previously, the edge spectra of non-Hermitian systems under OBCs are obtainable via the edge matrix [78]. Consider the $m$-band non-Hermitian Hamiltonian $H(\beta)=\sum_{i=-r}^{l}h_{i}\beta^{i}$ after the substitution $e^{ik}\rightarrow\beta$, the block element of the $mr\times mr$ edge matrix $M_{\text{matrix}}$ is defined as [78]:
@@
where $a,b=1,2,\cdots,r$, and the counterclockwise integral contour encloses the origin and the first $p$ solutions $\beta_{1}(E),\cdots,\beta_{p}(E)$ of $\det(H(\beta)-E)=0$. The isolated edge modes are determined by the relation $\det[M_{\text{edge}}]=0$ [85]. Compared to the edge matrix, the recurrence method is more efficient, as it restricts the edge-spectrum problem to the few zero-coefficient solutions of $d_{i}(E)=0$, while through the edge matrix, $\det[M_{\text{edge}}(E)]\neq 0$ a complicated and costly matrix equation, especially when $mr$ is large - has to be verified for all $E$ outside the OBC spectra through the edge matrix [78].
@@ IV.2 Non-Hermitian size-parity effect
where $i$ labels the unit cell and $\tau=A,B$ denotes the sublattices. $\delta_{\tau,A}$ indicates that the edge mode weighs exclusively on the $A$ sublattice. It is apparent from the decay constant $\kappa$ that such a localized edge mode exists only when $v<w$. Likewise, when and only when $v<w$, a right-localized edge mode with $E=-V$ emerges on the $B$ sublattices at the right open boundary.
@@ V Non-Hermitian Hofstadter model
where the NN hoppings $t_{x}$, $t_{x}^{\prime}$, and the onsite potential $V_{x}$ are in the Supplemental Materials [74] and generally complex-valued; therefore, the comparability with and similarity transformation to a Hermitian model no longer holds in such triangular-lattice Hofstadter butterfly models. Also, we consider a commensurate perpendicular magnetic field $B_{r}$, i.e., $B_{r}=2\pi(p/q)$ with coprime integers $p$ and $q$. Consequently, the Hamiltonian in Eq. (44) is $q$-periodic ($k=q$). However, numerical diagonalization of a non-Hermitian Hamiltonian is highly susceptible to accumulated errors and sensitive to numerical precisions, especially in large systems [47, 96, 13]. The non-Bloch band theory and GBZ method also face substantial obstacles and challenges for relatively large $q$, as the number of sublattices in the magnetic unit cell significantly elevates the computational cost. Hence, trustworthy calculations of energy spectra of Eq. (44) for large $N$ and $q$ are exceedingly time-consuming [47]. Through the recurrence method, on the other hand, we transform the eigenvalue problem of Eq. (44) to the root solutions of a series of $q$-th degree polynomials, which share the same coefficients except for the constant terms, as we have shown in Sec. III. Numerically, we can solve these polynomials with an iterative algorithm called the Durand-Kerner method [76, 77, 74], efficient and useful for simultaneously determining all complex roots of high-degree polynomials with a quadratic convergence rate. Therefore, the recurrence method avoids the numerical accuracy and instability issue in diagonalizing large non-Hermitian matrices and substantially reduces the computational complexity to $O(N)$ as long as $q\ll N$.
@@
For example, we examine the edge modes of the non-Hermitian Hofstadter models, i.e., Eq. (42) in the presence of the perpendicular magnetic field, especially their dispersions with respect to $k_{y}$ the momentum along the open boundaries. We show example results for $p/q=1/7$ in Fig. 6, which exhibit apparent edge modes (red), obtained from condition (ii) of the recurrence method, traversing the gaps between the bulk bands (black), mainly obtained from condition (i) of the recurrence method. Interestingly, compared with the Hermitian case in Fig. 6(a), although the non-reciprocity $\delta=0.5$ essentially deforms the bulk and edge spectra, the presence of the topological edge modes persists in the non-Hermitian Hofstadter models, as illustrated in Fig. 6(b). This is smoking-gun evidence that the corresponding topological phases generalize to non-Hermitian scenarios and, simultaneously, offers a clear-cut identification of the corresponding Chern numbers as $+2$  [^1] The separation between the lowest two bulk bands is unclear. Thus, we present their topological characterization as a whole. , $+1$, $-1$, $-1$, $-1$ of the bulk bands from the bottom to the top.

### v2

@@ I Introduction
Non-Hermiticity is a significant factor that can not be overlooked in both classical and quantum systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. In these systems, non-Hermiticity gives rise to various unconventional phenomena such as non-Hermitian skin effect [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], non-Bloch parity-time symmetry [31, 32, 33, 34], non-Hermitian topology [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], generalized Brillouin zone [47, 45, 48, 49, 50], and exceptional points [51, 52, 53, 54, 55, 56]. One intriguing property of non-Hermitian systems is the breakdown of spectral stability under different boundary conditions [57, 58, 59, 60, 61]. Specifically, it is widely believed that the energy levels of Hermitian systems with an open boundary condition (OBC) can be reproduced mainly by the Bloch band theory under periodic boundary conditions (PBCs), even if the translation symmetry is broken. In non-Hermitian systems, however, the energy spectra under OBC change drastically and collapse into open arcs encircled by the loop-shape periodic boundary spectra [62, 45, 21, 63]. It is widely known that these non-Hermitian OBC spectra can be obtained by the non-Bloch theory, which extends the Brillouin zone to the generalized Brillouin zone (GBZ) in non-Hermitian systems [45, 49, 48, 21]. Unfortunately, the determination of GBZ is somewhat challenging; therefore, it may be necessary to use the auxiliary GBZ technique [48], which still involves additional concepts and may not apply to complicated multi-band non-Hermitian systems, such as non-Hermitian Hofstadter butterflies [48]. Furthermore, non-Bloch theory primarily emphasizes the bulk OBC spectra relevant to the GBZ, and specialized tools for edge spectra of non-Hermitian systems, intimately connected with topological properties, are still largely lacking.
@@
To circumvent the limitations of the GBZs and the non-Bloch band theory, we propose in this paper an alternative, general, efficient, and accurate method based on the recurrence relations of the eigenvalue equations in non-Hermitian systems under OBC. First, we conclude that the characteristic polynomial $D_{N}(E)=\det(E-H_{N})$ of any non-Hermitian tight-binding Hamiltonian under OBC satisfies a recurrence relation with constant coefficients independent of the matrix size $N$. Then, we can represent $D_{N}(E)$ with the solutions $z_{i}(E)$ of the corresponding characteristic equation. The main result is that the OBC spectra include the energies such that (i) two maximal complex solutions $z_{i}$ and $z_{j}$ have the same modulus, or (ii) when representing $D_{N}$ by $z_{i}$, the maximal complex solution term is zero. We will argue that the above conditions are related to the bulk and edge parts of OBC spectra, respectively.
@@
As an application of the recurrence method, we derive a general expression for the bulk OBC spectra of non-Hermitian multi-band Hamiltonians with nearest-neighbor hoppings. This expression has the form of polynomial equations that allow us to analytically or numerically solve the OBC spectra without either diagonalization or determination of the GBZ. Thus, the recurrence method avoids typical accuracy problems in diagonalizing non-Hermitian matrices and significantly reduces the computational complexity. Furthermore, our formalism straightforwardly determines non-Hermitian Hamiltonians’ edge spectra, which are unavailable from the conventional non-Bloch band theory. Based on this framework, we can analyze the effects of the overall number of sites (e.g., the size-parity effect) [64, 65, 66] and the boundary perturbations [67, 68, 69, 70] on the edge spectra. Such bulk and edge spectra offers direct information on physical properties such as the PT-symmetry breaking, topological characters, and bulk-boundary correspondence of the target non-Hermitian systems under OBCs. For example, we study the non-Hermitian Hofstadter model in two dimensions as an interplay between the non-reciprocity and magnetic field [71]. Our results provide robust evidence for the complex semiclassical theory [72] and topological characterizations of non-Hermitian systems [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Generally, our work offers a novel technique and perspective for non-Hermitian systems’ OBC spectra and physics.
@@ II Recurrence Method in non-Hermitian systems
where $d_{i}(E)$ are determined by the initial conditions $D_{0}(E)=1,D_{k}(E),\cdots,D_{(s-1)k}(E)$ of the recurrence relation in Eq. (3), which are the characteristic polynomials in Eq. (2) corresponding to Hamiltonian matrices of size $0,k,2k,\cdots,(s-1)k$, respectively. The characteristic polynomial of the null matrix (size zero) is conventionally defined as $1$, while the others can be computed directly through the standard determinant calculations in general, which are relatively straightforward due to their (initially) limited sizes. The OBC spectrum in the thermodynamic limit $N\rightarrow\infty$ is comprised of values of $E$ satisfying either of the following two conditions [76, 75]:
@@
Note that condition (i) contains an infinite number of $E$ in a continuous spectrum of solutions and a similar form with the GBZ condition $|\beta_{p}(E)|=|\beta_{p+1}(E)|$ in the non-Bloch band theory, thus corresponding to the continuous bulk OBC spectra of the non-Hermitian systems. For example, the recurrence relation for the Hatano-Nelson model:
$H_{\text{HN}}=\sum_{i}[(t+\gamma)c_{i+1}^{\dagger}c_{i}+(t-\gamma)c_{i}^{\dagger}c_{i+1}],$ (6)
is:
$D_{N}^{\text{HN}}(E)=ED_{N-1}^{\text{HN}}(E)-(t^{2}-\gamma^{2})D_{N-2}^{\text{HN}}(E),$ (7)
whose characteristic equation:
$z^{2}=Ez-(t^{2}-\gamma^{2}),$ (8)
has a degree $s=2$. Condition (i) $|z_{1}|=|z_{2}|=\sqrt{t^{2}-\gamma^{2}}$ offers a continuous set of solutions and differs from the GBZ condition $|\beta_{1}|=|\beta_{2}|=\sqrt{(t+\gamma)/(t-\gamma)}$ only by a constant. As a result, the spectrum $E=2\sqrt{t^{2}-\gamma^{2}}\cos\theta$ is consistent with the non-Bloch band theory [15].
On the other hand, condition (ii), combined with the initial conditions of Eq. (5), gives several polynomial equations about $E$ and thus a discrete set of solutions. These polynomial equations are generally quite different from the non-polynomial form of condition (i) and generally insensitive of the system size $N$. Hence, we expect condition (ii) to contain the edge spectra of non-Hermitian systems under OBCs, which is beyond the scope of the non-Bloch band theory and the GBZ with only information of the bulk spectra [49, 48]. For example, the spectrum equation of the Hatano-Nelson model following condition (ii) and approach elaborated in subsequent sections is:
$t^{2}-\gamma^{2}=0,$ (9)
which yields zero solutions, thus edge modes, which is consistent with the absence of edge modes in the single-band Hatano-Nelson model and, more generally, condition (ii) for the edge spectra.
The recurrence method is remarkably useful for non-Hermitian models, especially models with large unit cells ($k\gg 1$) that are exceedingly complex from a non-Bloch theory perspective. As we will show, even for a large unit cell, the recurrence method gives an analytic expression for the OBC spectra of 1D non-Hermitian nearest-neighbor tight-binding models in the thermodynamic limit $N\rightarrow\infty$. Computationally, on the other hand, the recurrence method transfers the OBC spectrum problem to the problem of solving polynomials of degree $k$, which reduces the time complexity to $O(Nk^{2})$ with the Durand-Kerner algorithm [77, 75].
@@ III.1 Periodic nearest-neighbor-hopping non-Hermitian models
which is an analytical expression for the OBC spectra of multiple-band Hamiltonians in Eq. (10) through the recurrence method. Therefore, our formalism transforms the OBC spectra problems of Eq. (10) to a sequence of polynomial equations in Eq. (17). For a lattice period or sublattice size $k$ that is arbitrarily large, we can solve Eq. (17) numerically through an iterative method called the Durand-Kerner method [77, 75] with manageable cost $\sim O(Nk^{2})$, advantageous over the non-Bloch band theory.
@@ III.2 Non-Hermitian Rice-Mele model
where $w_{1/2}$ denote the intercell hoppings between the adjacent unit cells, $v_{1/2}$ denote the intracell hoppings between sites A and B in the same unit cell, and $V$ is a staggered onsite potential, as shown in Fig. 1(a). The $V=0$ case recovers the non-Hermitian SSH model in Eq. (19) with vanishing onsite potentials. The asymmetric intracell/intercell hoppings, e.g., $v_{1}=v-\gamma$ versus $v_{2}=v+\gamma$ with a finite $\gamma$, introduce non-Hermiticity into the model and can be realized in experiments by reservoir engineering in the ultra-cold atom systems and imaginary gauge field in coupled optical micro-ring resonators [36, 82, 83].
@@
Figure 1: We compare the OBC (red curves and dots) and PBC (blue curves) spectra of the non-Hermitian Rice-Mele model in Eq. (25) with two sublattices A and B and asymmetric hoppings $v_{1}\neq v_{2}^{*}$ or $w_{1}\neq w_{2}^{*}$, illustrated in (a). The OBC spectra are fully real when both $v_{1}v_{2},w_{1}w_{2}>0$, irrespective of $V$: (b) $v_{1}=2$, $v_{2}=1$, $w_{1}=1.2$, $w_{2}=0.8$, and $V=0.5$; (c) $v_{1}=0.8$, $v_{2}=0.2$, $w_{1}=1.2$, $w_{2}=0.8$, and $V=0.5$; (d) $v_{1}=0.8$, $v_{2}=0.2$, $w_{1}=-1$, $w_{2}=1$, and $V=0.5$. However, when both $v_{1}v_{2},w_{1}w_{2}<0$, a non-Bloch PT symmetry breaking may occur: (e-g) $v_{1}=-0.8$, $v_{2}=0.2$, $w_{1}=-1$, and $w_{2}=1$, thus the transition points are at $V_{c}=1.4$. The spectrum is complex for (d) and (e) $V=1.2$ at a critical point at (f) $V=1.4$ and becomes fully real for (g) $V=1.5$. Note that $|v_{1}v_{2}|>|w_{1}w_{2}|$ and thus no edge state exists in (b), while $|v_{1}v_{2}|<|w_{1}w_{2}|$ and two isolated edge states appear at $E=\pm V$ (red dots) in (c-g).
@@ IV.1 Edge spectra by recurrence method
Previously, the edge spectra of non-Hermitian systems under OBCs are obtainable via the edge matrix [78]. Consider the $m$-band non-Hermitian Hamiltonian $H(\beta)=\sum_{i=-r}^{l}h_{i}\beta^{i}$ after the substitution $e^{ik}\rightarrow\beta$, the block element of the $mr\times mr$ edge matrix $M_{\text{edge}}$ is defined as [78]:
@@
where $a,b=1,2,\cdots,r$, and the counterclockwise integral contour encloses the origin and the first $p$ solutions $\beta_{1}(E),\cdots,\beta_{p}(E)$ of $\det(H(\beta)-E)=0$. The isolated edge modes are determined by the relation $\det M_{\text{edge}}=0$ [85]. Compared to the edge matrix, the recurrence method is more efficient, as it restricts the edge-spectrum problem to the few zero-coefficient solutions of $d_{i}(E)=0$, while through the edge matrix, $\det M_{\text{edge}}(E)\neq 0$ a complicated and costly matrix equation, especially when $mr$ is large - has to be verified for all $E$ outside the OBC spectra through the edge matrix [78].
@@ IV.2 Non-Hermitian size-parity effect
where $i$ labels the unit cell and $\tau=A,B$ denotes the sublattices. $\delta_{\tau,A}$ indicates that the edge mode weighs exclusively on the $A$ sublattice. It is apparent from the decay constant $\kappa$ that such a localized edge mode exists only when $v<w$. Likewise, when and only when $v>w$, a right-localized edge mode with $E=-V$ emerges on the $B$ sublattices at the right open boundary.
@@ V Non-Hermitian Hofstadter model
where the NN hoppings $t_{x}$, $t_{x}^{\prime}$, and the onsite potential $V_{x}$ are in the Supplemental Materials [75] and generally complex-valued; therefore, the comparability with and similarity transformation to a Hermitian model no longer holds in such triangular-lattice Hofstadter butterfly models. Also, we consider a commensurate perpendicular magnetic field $B_{r}$, i.e., $B_{r}=2\pi(p/q)$ with coprime integers $p$ and $q$. Consequently, the Hamiltonian in Eq. (48) is $q$-periodic ($k=q$). However, numerical diagonalization of a non-Hermitian Hamiltonian is highly susceptible to accumulated errors and sensitive to numerical precisions, especially in large systems [48, 96, 13]. The non-Bloch band theory and GBZ method also face substantial obstacles and challenges for relatively large $q$, as the number of sublattices in the magnetic unit cell significantly elevates the computational cost. Hence, trustworthy calculations of energy spectra of Eq. (48) for large $N$ and $q$ are exceedingly time-consuming [48]. Through the recurrence method, on the other hand, we transform the eigenvalue problem of Eq. (48) to the root solutions of a series of $q$-th degree polynomials, which share the same coefficients except for the constant terms, as we have shown in Sec. III. Numerically, we can solve these polynomials with an iterative algorithm called the Durand-Kerner method [77, 75], efficient and useful for simultaneously determining all complex roots of high-degree polynomials with a quadratic convergence rate. Therefore, the recurrence method avoids the numerical accuracy and instability issue in diagonalizing large non-Hermitian matrices and substantially reduces the computational complexity to $O(N)$ as long as $q\ll N$.
@@
For example, we examine the edge modes of the non-Hermitian Hofstadter models, i.e., Eq. (46) in the presence of the perpendicular magnetic field, especially their dispersions with respect to $k_{y}$ the momentum along the open boundaries. We show example results for $p/q=1/7$ in Fig. 6, which exhibit apparent edge modes (red), obtained from condition (ii) of the recurrence method, traversing the gaps between the bulk bands (black), mainly obtained from condition (i) of the recurrence method. Interestingly, compared with the Hermitian case in Fig. 6(a), although the non-reciprocity $\delta=0.5$ essentially deforms the bulk and edge spectra, the presence of the topological edge modes persists in the non-Hermitian Hofstadter models, as illustrated in Fig. 6(b). This is smoking-gun evidence that the corresponding topological phases generalize to non-Hermitian scenarios and, simultaneously, offers a clear-cut identification of the corresponding Chern numbers as $+2$  [^1] The separation between the lowest two bulk bands is unclear. Thus, we present their topological characterization as a whole. , $+1$, $-1$, $-1$, $-1$ of the bulk bands from the bottom to the top, as conventionally in Hermitian systems, the (signed) number of chiral edge modes traversing a gap determines the total topological invariants across all bands below the gap.